Please explain why the graph of (x-h)^2 = 4p(y-k) has a y-intercept of
(0, h^2/4p +K). Thanks.

Answer by Nate(3500)

(x-h)^2 = 4p(y-k)
(0-h)^2 = 4p(y-k)
h^2 = 4p(y-k)
h^2/4p = y-k
h^2/4p+k = y